OpenStudy (anonymous):

Can someone please explain me how u can get x=5, x=-1 after seeing this (x)=(x−5)(x+1)? xD

Nnesha (nnesha):

set both parentheses equal to 0 x-5 =0 x+1 =0 solve for x

OpenStudy (anonymous):

is that the quadratic

OpenStudy (anonymous):

no never mind sorry

Nnesha (nnesha):

yep those 2 parentheses are factors of quadratic equation \[\huge\rm Ax^2+Bx+C=0\]